The search "A kite is on a 150 metre string that a boy holds 1metre above the ground. The horizontal distance of the kite from the boy is 80metres. How high is the kite above the ground?
Answer by Alan3354(69443)

A kite is on a 150 metre string that a boy holds 1metre above the ground. The horizontal distance of the kite from the boy is 80metres. How high is the kite above the ground?
----------
A = angle of the string with the ground
cos(A) = 80/150
side opposite A = 150*sin(A)
height = 150*sin(A) + 1